Mixing 100% antifreeze with 40% antifreeze to get 20 quarts of 50% antifreeze
Answer by checkley71(8403)

X+.4(20-X)=.5*20
X+8-.4X=10
.6X=10-8
.6X=2
X=2/.6
X=3.333 QTS OF PURE ANTIFREEZE IS NEEDED TO GET A 50% MIX.
PROOF
3.333+.4(20-3.333)=.5*20
3.333+.4*16.667=10
3.333+6.667=0
10=10